OpenStudy (anonymous):

What is the period of the graph of y = 2tan(3*pi*x + 4)?

OpenStudy (anonymous):

\[\frac{2\pi}{b}\]

OpenStudy (anonymous):

in this case \[b=3\pi\]

OpenStudy (anonymous):

so you get \[\frac{2}{3}\]

OpenStudy (anonymous):

whooops!

OpenStudy (anonymous):

this is tangent!

OpenStudy (amistre64):

someone forgot to adjust their bifocals lol

OpenStudy (anonymous):

i was just about to say that haha. period is Pi for tan(x)

OpenStudy (anonymous):

sorry. period of tangent is pi so i lied

OpenStudy (anonymous):

ok ok ok

OpenStudy (anonymous):

1/3 so sue me

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Hold on, what exactly did you do satellite73?

OpenStudy (anonymous):

i made a mistake. it is not \[\frac{2\pi}{b}\]

OpenStudy (anonymous):

it is \[\frac{\pi}{b}\]

OpenStudy (anonymous):

b is the coefficient of the x which in your case is \[3\pi\]

OpenStudy (anonymous):

so i just computed \[\frac{\pi}{3\pi}\] and got \[\frac{1}{3}\]

OpenStudy (anonymous):

as the pi's cancel

OpenStudy (anonymous):

sorry i screwed up on the first try

OpenStudy (anonymous):

Haha, np